OpenStudy (anonymous):

Help please? Write the standard form of a line that contains a slope of 2/3 and passes through the point (1,1).

OpenStudy (anonymous):

actually, lets make it even easier. standard form is simply mx + y = b

OpenStudy (anonymous):

That's what's confusing me. How am I supposed to use the points here?

OpenStudy (anonymous):

points are of the form (x,y), we can simply sub in the equation.

OpenStudy (anonymous):

we are given m = 2/3. so our equation now is: (2/3)x + y = b.

OpenStudy (anonymous):

why dont you try subbing in x = 1, and y = 1 into the equation and solve for b? that will be your solution.

OpenStudy (anonymous):

So you simply just plug in those two points for x and y turning the equation into y=2/3x+1?

OpenStudy (anonymous):

no sorry, try subbing into the equation we found above: -(2/3)x + y = b. (forgot to include the - last time! sorry) try finding b.

OpenStudy (anonymous):

So instead -(2/3)1+1=b , would you then multiply -(2/3) into 1 making it positive and turning the equation into 2/3+1=b?

OpenStudy (anonymous):

not quite. remember a negative (-2/3 in this case) and positive (1) = negative, so we would get back -2/3. simply sub in (x = 1 and y = 1) -(2/3)1+1=b -2/3 + 1 = b so what is b?

OpenStudy (anonymous):

Is it one?

OpenStudy (anonymous):

what is (-2/3) + 1, try it on your calculator. it is not 1.

OpenStudy (anonymous):

It's 1/3

OpenStudy (nottim):

Yup.